OpenStudy (anonymous):

64 is 44 times the difference between Sarah's age, a, and 44. Assume Sarah is older than 44. Write an equation to determine Sarah's age (a). Find Sarah's age.

OpenStudy (anonymous):

@jasmineETTA

OpenStudy (anonymous):

@ganeshie8

OpenStudy (majesty69):

64/44=a x+44=sarah's age

OpenStudy (majesty69):

Are there options?

OpenStudy (anonymous):

thank you your@majesty69

OpenStudy (anonymous):

nope

OpenStudy (majesty69):

no problem x'D

OpenStudy (anonymous):

what does tht mean

OpenStudy (majesty69):

what

OpenStudy (anonymous):

XD?

OpenStudy (anonymous):

ohh hold up it also says what is her age?

OpenStudy (majesty69):

"x'D" is kinda of like an "lmfao" face...like laughing super hard....lol idk xD Well just follow that equation and you will find it

OpenStudy (anonymous):

wait whats the equation again i got mixed up

OpenStudy (majesty69):

64/44=a x+44=sarah's age

OpenStudy (anonymous):

i see u @jasmineETTA do u know what it is (on the viewing list

OpenStudy (jasmineetta):

i'm not really good with math

OpenStudy (anonymous):

ohhh yea 4got

OpenStudy (anonymous):

but i got 45.4545454545 is tht it your@majesty69

OpenStudy (majesty69):

yup so she's about 45 1/2

OpenStudy (anonymous):

ohhh ok thx

OpenStudy (jasmineetta):

i'm not really good with math